# KEN KANGDA WREN1

### v7

@@ Introduction
Let our model-risk $B$ have two outcomes, $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary (’model-risk free’). Such $B$-sure processes represent ’routines’ that by classical RE principles are well-modelled; only change creates model-risk (e.g. new CEO, R&D results, macroeconomic, regulatory or geopolitical events). Non-binary situations that can be handled within the binary framework emerge quite naturally, as will be demonstrated in an application addressing typical questions encountered in practice.
@@
Some model-risks of this type may be replicated/hedged or diversified. Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
Our model-risk shall be considered one-off. How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time. This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated. Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g. Hansen et al. (2006)). Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e. while the given model-risk remains active and relevant).
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the objectively true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
Lastly, data on model-risk (e.g. weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data on ’routine’ risks (e.g. biannual dividend). Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
The setup as outlined above makes model-risk a prominent feature of asset-pricing, a major source of excess value, and so a focus of market effort and activities.
From the Perspectives of Standard Model-Risk Studies
@@
Finally, existing theories require ’total risk’ as a whole as inputs, model and non-model risks together, aggregated across all relevant assets in a given market. They are unwieldy for bottom-up (individual-asset) or cross-sectional enquiries, especially when the model-risk(s) can be micro-economic or heterogeneous in origin. Our method, with its isolation of model-risk pricing and RNE formulation, is more flexible and convenient for such applications.
@@
Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
We find that the RNE formulations of asset-pricing under model-risk are constrained by economic, informational, necessities (Definition 1 and Lemma 1) and so have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2). One implication is a potential link between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves reveal and separate ex-ante risk-premia from bias under model-uncertainty (Section 3.4.1&3.4.2).
@@ 1.1 The Usual-Case Asset and Information Structure
For clarity, consider an usual-case asset: it has just one, bullet, payoff, $Y\in\mathbb{R}$, at some horizon or maturity $T$ much longer than unit-time, so $1\ll T<\infty$; let value and price be in logarithm, and time- and risk-free discount exogenous and nil (in a suitable numéraire). Further:
@@
    Let $Y$ be revealed continuously by a public record $\{Z_{t}\}$ of realised increments (decrements) to some known base, set to $\log 1=0$, where each $Z_{t}$ is the sum of such firmed-up changes to date; by setup $Y\equiv Z_{T}$. Let $\{Z_{t}\}$ be a diffusion, based on some standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to outcome $\{B=b\}$, $b\in\mathcal{B}:=\{+,-\}$,
    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable. The underlying $B$-sure space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$. The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g. Peskir and Shirayev (2006)) concern processes of fixed drift and volatility. Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities. For more background see Appendix A..
@@
    The economic impact of model-risk $B$ thus derives from drift difference $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$; let it be consistent: $Y^{\Delta}(t):=\int_{t}^{T}r^{Z\Delta}_{s}ds>0$ almost surely, at any $t<T$. This is intuitive, also necessary for viable and economically valid risk-pricing, as will be seen (Remark 8).
@@
    There may be a purely $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$ that is $B$-conditionally independent but otherwise similar to $\{Z_{t}(B)\}$, based on a standard Wiener process $\{w^{D}_{t}\}$,
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}.$ (2)
    It by definition is price-irrelevant in the absence of model-risk (when $B$-sure). Combined, total dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under the product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, with $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$.
@@
    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible firmed-up value, $\{Z_{t}\}$, exist/matter. By considering data exclusively relevant to model-risk (irrelevant once model-sure), we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate asset valuation and hence traders’ cognitive and trading activities, while non-model, routine, risks remain, independently evolving or hardly changing.
  4. 4.
[…]
@@
    Tests based on likelihood-ratios (LR) $L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}$, or log-LR $l^{+/-}_{t}:=\log L^{+/-}_{t}$, $\forall t<T$, are optimal in the sense of Neyman-Pearson Lemma. Model-risk $B$ resolves iff. $\exists T_{\mathcal{B}}\leq\infty$ such that $\lim_{t\to T_{\mathcal{B}}}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T_{\mathcal{B}}}\perp\mathbf{W}_{-}|_{T_{\mathcal{B}}}$). Further, such inferential tests are performed under the equivalent martingale measure of one of the processes (e.g that of $B=$-). Henceforth let $\mathbf{W}_{-}$ be this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
@@
  7. 7.
    Given $Z_{t}=z_{t}$, $t<T$, the $B$-sure expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]\equiv{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]$ reads ((1-2)):
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
    where $y_{B}(T-t,t)-y_{B}(T-t-t^{\prime},t+t^{\prime})=\mathbf{E}_{B}[\int_{t}^{t+t^{\prime}}dZ_{s}(B)]$, $\forall t^{\prime}\leq T-t$, and so $y_{B}(T-t,t)=\int_{t}^{T}{r}^{Z}_{B,s}ds$, as by definition $\{Y^{B}_{t}\}$ is a martingale and $y_{B}(0,T)\equiv 0$. Note also that this relates to model-risk impact $Y^{\Delta}(t)$ via $Y^{\Delta}(t)\equiv Y^{+}_{t}-Y^{-}_{t}\equiv{y}_{+}(T-t,t)-{y}_{-}(T-t,t)$.
@@ 1.2 The FTAP Baseline under Model-Risk
Viability by no means ensures economic validity (e.g. any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-informative data); indeed, any pricing that is viable to any member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class. In this sense, ’anything goes’ under FTAP-viability. What characterises the processes and the RNE measures that do correspond to plausible economic theories?
For answers, consider risk-premium, the gap between ex-ante pricing and expectation, non-negative under economic asset-pricing: at $\forall t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0.$ (7)
@@
where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ accounts for $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$. The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity. It well-defines a RNE price-of-$B$-risk:

### v8

@@ Introduction
Let our model-risk $B$ have two outcomes, $\{B=b\}$ or $\{B=\overline{b}\not=b\}$, and our $B$-sure models, free of model-risk. The $B$-sure processes represent ’routines’, which by classical RE principles must be well-modelled. Only change may create model-risk (e.g. new CEO, R&D results, macroeconomic, regulatory or geopolitical events). Non-binary situations that can be handled within the binary framework emerge quite naturally, as will be demonstrated in an application addressing typical questions encountered in practice.
@@
Some risks of this type may be diversified away, or replicable and so hedged out. Our study concerns only the economically priced: any deemed worthy of being given at each moment a risk-premium via some recursive optimisation scheme with said risk as an input.
Model-risk $B$ is to be treated as one-off. How it ’flares up’ and the pricing of ’flare risk’ are implicit; that is, we study episodic ’model crises’, one at a time. Such a backdrop puts the core issues in sharp relief and is a good approximation of both real-world settings and the favourite of the literature, Markov hidden regime-switching that is rare. Let horizon or maturity be long enough for meaningful model-risk inference to develop, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such periods.
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the objectively true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual strategy and ours.
Our main point of departure from existing treatments is the separation of model-risk from the rest, the non-model risks, in terms of its inference and discounting, to arrive at a tractable characterisation of RNE asset-pricing under model-risk despite path-dependency.
The separation is reflected also in our information structure, through the potential presence of data exclusive to model-risk, as opposed to the standard one, which is generated by the irreducible inputs of asset-worth alone (e.g. dividends). It mimics the common scenario where data about model-risk (e.g. economic indicators) are far more varied and frequent than the realisations of irreducibles (e.g. earnings). It also enables and enriches the economic analysis of uncertainties with sure pre-horizon resolution. Such uncertainties, although common in practice, fall outside the remit of current theories, which require model or parameter risks to stay unresolved, or as in Hansen-Sargent’s robust-control, resolved at infinity (e.g. Hansen et al. (2006)), since risks surely resolvable pre-horizon in a standard information structure are replicable/hedgeable and so priced by arbitrage. Our setup accommodates unreplicable model-risks that are however sure to resolve continuously before horizon or maturity.
The Perspectives of Standard Model-Risk Studies
@@
Section 1 sets out the asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
The RNE formulations of asset-pricing under model-risk are shown to be constrained by economic, informational, necessities (Definition 1 and Lemma 1) to have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2). One implication is a potential link, given model-uncertainty, between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves can identify and separate ex-ante risk-premia from bias (Section 3.4.1&3.4.2).
@@ 1.1 The Usual-Case Asset and Information Structure
Consider an usual-case asset, one with a sole payoff $Y\in\mathbb{R}$, in logarithmic value, at a far off horizon $T$, $1\ll T<\infty$, with time- and risk-free discount set to nil (in a suitable numéraire):
@@
    Let $Y$ be revealed continuously by a public record $\{Z_{t}\}$ of realised incremental changes to a known base level, set to $\log 1=0$, where each $Z_{t}$ is the sum of firmed-up changes to date, with $Z_{T}\equiv Y$ by definition. Let $\{Z_{t}\}$ be a diffusion, based on a standard Wiener process $\{w^{Z}_{t}\}$, generating standard Wiener space $\mathcal{S}^{Z}$, measure $w^{Z}$ and natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to outcome $\{B=b\}$, $b\in\mathcal{B}:=\{+,-\}$, where its $B$-sure drift $\{r^{Z}_{B,t}\}$ and $B$-independent volatility $\{\sigma^{Z}_{t}\}$ are $\{\mathcal{F}^{Z}_{t}\}$-predictable:
    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in[0,\infty).$ (1)
    The Wiener space $(\mathcal{S}^{Z},W^{Z}_{B})$ of the $B$-sure process $\{Z_{t}(B)\}$, with $dW^{Z}_{B,t}=dZ_{t}(B)$, provides the state space and measure $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ of $B$-sure payoffs via restrictions to $T$. Model-risk inference amounts to the standard sequential testing of Wiener processes (Peskir and Shirayev (2006)), assuming ${W}^{Z}_{+}\perp{W}^{Z}_{-}$ and ${W}^{Z}_{+}|_{t}\sim{W}^{Z}_{-}|_{t}$, $\forall t<\infty$, so that data $\{Z_{t}\}$ allow $B$-resolution at $T^{Z}_{\mathcal{B}}:=\inf\big{\{}t|\mathcal{F}^{Z}_{t}\ni\{B=b\}\big{\}}=\infty$ almost surely [^2] Text-book cases concern constant drifts and volatility, from which tests with $\{\mathcal{F}^{Z}_{t}\}$-predictable parameters derive via an absolutely continuous clock-change (see e.g. Kallsen (2006) and also Appendix A)..
@@
    Then, given $Z_{t}=z_{t}$, $t<T$, the $B$-sure payoff expectation process $\{Y^{B}_{t}\}:=\{{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]\}$, a well-known martingale, takes the following form:
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=:z_{t}+y_{B}(T-t,t);$ (2)
    note $y_{B}(0,T)\equiv 0$ and $y_{B}(T-t,t)=\mathbf{E}_{B}[\int_{t}^{T}dZ_{t}]=\mathbf{E}[\int_{t}^{T}{r}^{Z}_{B,s}ds]$, where $\mathbf{E}[\cdot|\cdot]$ denotes expectations under the standard Wiener measure $w^{Z}$.
@@
    Define model-risk impact $Y^{\Delta}(t):=Y^{+}_{t}-Y^{-}_{t}=\mathbf{E}[\int_{t}^{T}r^{Z\Delta}_{s}ds]$, with $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$. Let it be consistent: $Y^{\Delta}(t)>0$, $\forall t<T$. This is intuitive, and necessary, as it turns out, for viable and economically valid risk-pricing (Remark 9).
  4. 4.
    There is a purely $B$-informative real-vector dataflow $\{{D}_{t}\}$ that is $B$-conditionally independent to $\{Z_{t}\}$ (so price-irrelevant if given $B$) based on standard Wiener process $\{w^{D}_{t}\}$,
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}=:dW^{D}_{B,t},\text{ }t\in[0,T^{D}_{\mathcal{B}}),$ (3)
    where $T^{D}_{\mathcal{B}}:=\inf\big{\{}t|\mathcal{F}^{D}_{t}\ni\{B=b\}\big{\}}\leq\infty$, that is, ${W}^{D}_{+}\perp{W}^{D}_{-}$ and ${W}^{D}_{+}|_{t}\sim{W}^{D}_{-}|_{t}$, $\forall t<T^{D}_{\mathcal{B}}$. As such, data $\{{D}_{t}\}$ may resolve $B$ in finite time (see Appendix A for the construction of such processes from standard diffusion). Model-risk $B$ is thus active on $0,T^{D}_{\mathcal{B}}\wedge T)$.
@@
    Existing studies assume, de facto, that only data feeding directly into tangible, firmed-up, values exist/matter. By adding data exclusive to model-risk and irrelevant if model-sure, we may tackle a frequent challenge facing partitioners: a market unsettled by a bout of model-risk whose economics dominate asset value and so market activities, while routine risks remain, independently evolving, in the background.
@@
    Combined, asset data $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ have $B$-sure state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$. Overall model-risk inference takes place on filtered space ($)\big{(}\mathcal{S}^{T}_{\mathcal{B}}:=\mathcal{B}\times\mathcal{S}^{T}\ni\omega_{B}:=(B,\omega),\text{ }\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\},\text{ }\mathbf{F}_{T};\text{ }{\pi}^{B(\cdot)}_{0}\mathbf{W}^{T}_{B(\cdot)}$($\cdot$$\big{)}$, where ${\pi}^{B(\omega_{B})}_{0}$ is the unconditional belief in outcome $B(\omega_{B}):=B$. Then, by Item-1&4, model-risk $B$ is active on $0,T_{\mathcal{B}}\wedge T)$, with $T_{\mathcal{B}}:=\inf\big{\{}t|\mathbf{F}_{t}\ni\{B=b\}\big{\}}={T}^{D}_{\mathcal{B}}$.
@@
    Tests based on likelihood-ratios (LR) $\{L^{+/-}_{t}\}:=\{\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}\}$, or log-LR $\{l^{+/-}_{t}\}:=\{\log L^{+/-}_{t}\}$, are optimal in the sense of Neyman-Pearson Lemma. Model-risk $B$ resolves iff. $\exists T_{\mathcal{B}}\leq\infty$ such that $\lim_{t\to T_{\mathcal{B}}}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T_{\mathcal{B}}}\perp\mathbf{W}_{-}|_{T_{\mathcal{B}}}$). Further, such inferential tests are performed under the equivalent martingale measure of one of the processes (e.g that of $B=$-). Henceforth let $\mathbf{W}_{-}$ be this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
  7. 7.
@@ 1.2 The FTAP Baseline under Model-Risk
Remark 4.
By setup, $\hat{S}^{B}_{t}=\hat{\mathbf{E}}_{B}\big{[}Y\big{|}\mathcal{F}^{Z}_{t}\big{]}=z_{t}+\hat{y}_{B}(T-t,t)$ given $Z_{t}=z_{t}$, as in (2). Thus, $B$-risk, which can be measurable (resolvable) to $\mathbf{F}_{T}$ but not $\mathcal{F}^{Z}_{T}$ (Item-1), is not measurable to $\mathcal{F}^{S}_{T}$, the price-generated $\sigma$-algebra via (5). It is not replicable and cannot be priced by arbitrage.
Viable prices do not mean economically valid prices. Indeed, any pricing that is viable to any one member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is then viable to all in the class; ’anything goes’ in this sense. What characterise processes and RNE measures that stem from valid economic theories? For answers, consider risk-premia $\{RP_{t}\}$, the gap between pricing and expectation, which should be non-negative, at least ex-ante: given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0\text{, }\forall t<T.$ (7)
@@
where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ discounts $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$. The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}{\pi}^{-}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity. It well-defines a RNE price-of-$B$-risk process $\{k_{t}^{\hat{\pi}}\}$: $\forall t<T$,